A transmission cable is carrying a current I = 22
Answer #1

8.-

The force is given by:

F = qv \times B

Where q is charge, v is speed and B is magnetic field. Also, the magnetic field produced by the wire is:

B = \frac{\mu _{0}i}{2\pi r}

Where i is current, miu is permeability of free space and r is distance. So, we first calculate the magnetic field:

B = \frac{(4\pi x10^{-7}Tm/A)(22A)}{2\pi (0.15m)}

B = 2.93x10^{-5}T

Now we can calculate the force:

F = qvB

F = (15x10^{-6}C)(55m/s)(2.93x10^{-5}T)

F = 2.42x10^{-8}N

ANSWER is A.

9.-

Using Snell's Law:

n = \frac{c}{v}

v = \frac{c}{n}

v = \frac{3x10^{8}m/s}{2.42}

v = 1.24x10^{8}m/s

Now, with the classic movement equation:

v = \frac{x}{t}

t = \frac{x}{v}

t = \frac{0.0026m}{1.24x10^{8}m/s}

ANSWER t = 2.1x10^{-11}s

I hope it helps!!